OpenStudy (unklerhaukus):

Calculate the expectation value of x squared \[\langle x^2\rangle \] \[\Psi(x,t) =\sqrt[4]{\frac{2am}{\pi\hbar}}e^{-a\left[\frac{mx^{2}}{\hbar}+it\right]}\]

OpenStudy (unklerhaukus):

\[\langle x^2 \rangle =\int\limits_{-\infty}^{\infty}x^2|\Psi|^2~\text d x\]

OpenStudy (unklerhaukus):

\[=\int\limits_{-\infty}^{\infty}x^2A^2e^{ -2a\left[\frac{mx^{2}}{\hbar}\right]}~\text d x\] \[=2\sqrt{\frac{2am}{\pi\hbar}}\int\limits_{0}^{\infty}x^2e^{ -2a\left[\frac{mx^{2}}{\hbar}\right]}~\text d x\] \[\text{let } \nu=\frac{2am}{\hbar}\] \[=2\sqrt{\frac{\nu}{\pi}}\int\limits_{0}^{\infty}x^2e^{ -\nu x^2}~\text d x\]

OpenStudy (unklerhaukus):

\(\int p \text dq=\left.pq\right|-\int q \text dp\) \(p=x^2\qquad\text dq= e^{-\nu x^2}\) \(\text dp=2x\text dx\qquad q= \frac{e^{-\nu x^2}}{-2x\nu}\) \[=2\sqrt{\frac{\nu}{\pi}}\left[\left.x^2\frac{e^{-\nu x^2}}{-2x\nu}\right|_{0}^{\infty}-\int\limits_{0}^{\infty}\frac{e^{-\nu x^2}}{-2x\nu}2x\text dx\right]\] \[=2\sqrt{\frac{\nu}{\pi}}\left[\left.x\frac{e^{-\nu x^2}}{-2\nu}\right|_{0}^{\infty}+\frac{1}{\nu}\int\limits_{0}^{\infty}{e^{-\nu x^2}}\text dx\right]\]

OpenStudy (unklerhaukus):

\[=\frac{2}{\nu}\sqrt{\frac{\nu}{\pi}}\int\limits_{0}^{\infty}{e^{-\nu x^2}}\text dx\] \[=\frac{2}{-2\nu^2 x}\sqrt{\frac{\nu}{\pi}}\left.{e^{-\nu x^2}}\right|_{0}^{\infty}\] \[=-\frac{1}{ x}\frac{\nu^{-3/2}}{\sqrt{\pi}}\left.{e^{-\nu x^2}}\right|_{0}^{\infty}\]

OpenStudy (unklerhaukus):

?

OpenStudy (phi):

your p should be x and dq x exp(stuff) dx

OpenStudy (anonymous):

i thought\[\int\limits_{0}^{\infty}e ^{-x ^{2}}d x=\frac{ \sqrt{\pi} }{ 2 } \] ???

OpenStudy (unklerhaukus):

but why @phi? good point @jeanlouie , what was i thinking

OpenStudy (phi):

you integrated dq and got a bogus result you want to integrate exp(u) du u in this case is x^2 (ignore constants) and du is 2 x dx so you need an x dx to integrate exp(x^2)

OpenStudy (unklerhaukus):

\[=\frac{2}{\nu}\sqrt{\frac{\nu}{\pi}}\int\limits_{0}^{\infty}{e^{-\nu x^2}}\text dx\]\[=\frac{2}{\nu}\sqrt{\frac{\nu}{\pi}}\sqrt{\frac{\pi}{4\nu}}\]\[=\frac 1 \nu\]\[=\frac1{{2am}/{\hbar}}\]\[=\frac{\hbar}{2am}\]

OpenStudy (unklerhaukus):

the answer in the back of my book says im off by a factor of two

OpenStudy (cruffo):

I gotta agree with @phi You would need chain rule to take the derivative of q.

OpenStudy (cruffo):

Sorry, not chain rule, quotient rule.

OpenStudy (unklerhaukus):

hmm

OpenStudy (unklerhaukus):

error function?

OpenStudy (cruffo):

What if you tried this:

OpenStudy (cruffo):

Kinda mess but:

OpenStudy (unklerhaukus):

\[=2\sqrt{\frac{\nu}{\pi}}\int\limits_{0}^{\infty}x^2e^{ -\nu x^2}~\text d x\]\(\int p \text dq=\left.pq\right|-\int q \text dp\) \(p=x\qquad\text dq= xe^{-\nu x^2}\text dx\) \(\text dp=\text dx\qquad q= \frac{e^{-\nu x^2}}{-2\nu}\) \[=2\sqrt{\frac{\nu}{\pi}}\left[\left.x\frac{e^{-\nu x^2}}{-2\nu}\right|_{0}^{\infty}+\frac{1}{2\nu}\int\limits_{0}^{\infty}{e^{-\nu x^2}}\text dx\right]\]\[=\frac{2{}\sqrt{\frac{\nu}{\pi}}}{2\nu}\int\limits_{0}^{\infty}{e^{-\nu x^2}}\text dx\]\[=\frac{1}{\nu}\sqrt{\frac{\nu}{\pi}}\int\limits_{0}^{\infty}{e^{-\nu x^2}}\text dx\]\[=\frac{1}{\nu}\sqrt{\frac{\nu}{\pi}}\sqrt{\frac{\pi}{4\nu}}\]\[=\frac 1 {2\nu}\]\[=\frac1{2\times2a[\frac{m}{\hbar}]}\]\[=\frac{\hbar}{4am}\]\[\Large\color{red}\checkmark\]

OpenStudy (cruffo):

Cool. Although I have to say I didn't understand a single bit of the first part of the problem!

OpenStudy (unklerhaukus):

thankyou to phi, cruffo, jeanlouie

OpenStudy (unklerhaukus):

what bit didn't you understand @cruffo?

OpenStudy (cruffo):

Well, I know what it means to find the expected value of a distribution. It's the first 3 lines of the discussion that I'm not sure what happened.

OpenStudy (cruffo):

Sorry, not of "of a distribution" should have been "of a random variable".

OpenStudy (unklerhaukus):

oh i think i used some information from previous parts of the question

OpenStudy (cruffo):

Ahhh.. Had to prove something about the Schrödinger equation in spherical coordinates in DE. What class is if for btw?

OpenStudy (unklerhaukus):

QM

OpenStudy (cruffo):

figures. Good luck!

OpenStudy (unklerhaukus):

\[\langle p \rangle = \int\limits_{-\infty}^{\infty}\frac{\hbar}{i} \frac{\partial}{\partial x} |\Psi|^2~\text d x\]\[=\int\limits_{-\infty}^{\infty}\frac{\hbar}{i} \frac{\partial}{\partial x} \sqrt{\frac{\nu}{\pi}}e^{ -\nu x^2}\text dx\]\[= \frac{\hbar}{i} \sqrt{\frac{\nu}{\pi}}\int\limits_{-\infty}^{\infty}\frac{\partial}{\partial x}e^{ -\nu x^2}\text dx\]\[= \frac{\hbar}{i} \sqrt{\frac{\nu}{\pi}}\int\limits_{-\infty}^{\infty}-2\nu xe^{ -\nu x^2}\text dx\]\[= -2\nu \frac{\hbar}{i} \sqrt{\frac{\nu}{\pi}}\int\limits_{-\infty}^{\infty}xe^{ -\nu x^2}\text dx\]\[=0 \qquad\text{(odd function)}\]